(Bond valuation) Flora Co.s bonds, maturing in 15 years, pay 12 percent interest on a $1,000 face value. However, interest is paid semiannually. If your required rate of return is 15 percent, what is the value of the bond? How would your answer change if the interest were paid annually? a. If the interest is paid semiannually, the value of the bond is $ (Round to the nearest cent.)please show work. Thank you!
Answer #1

Answer (a):

Bond face value = $1,000

Interest rate = 12%

Interest paid semiannually

Semiannual coupon payment = $1,000 * 12%/2 = $60

Time to maturity = 15 years

Number of semiannual periods = 15 * 2 = 30

Required rate of return = 15%

Semiannual required rate of return = 15%/2 = 7.5%

To get current value of bond we will use PV function of excel:

PV (rate, nper, pmt, fv, type)

PV (7.5%, 30, -60, -1000, 0)

=$822.84

If the interest is paid semiannually, the value of bond is $822.84

Answer (a) 2nd Part:

If interest is paid annually,

Current value =

PV (15%, 15, -120, -1000, 0)

= $824.58

Hence if interest is paid annually the current value of bond increases from $822.84 to $824.58